A student wearing frictionless roller skates on a horizontal surface is pushed by a friend with a constant force of 50.0 N. How FAR must the student be pushed, starting from rest, so that her final kinetic energy is 350 J?
1
Expert's answer
2020-03-30T07:50:53-0400

The change of energy is equal to work done

"K_f-K_i=W=Fd"

Hence, the distance

"d=\frac{K_f-K_i}{F}=\frac{350-0}{50.0}=7\:\rm m"